# A globally convergent gradient method with momentum

### v2

@@ Abstract
In this work, we consider smooth unconstrained optimization problems and we deal with the class of gradient methods with momentum, i.e., descent algorithms where the search direction is defined as a linear combination of the current gradient and the preceding search direction. This family of algorithms includes nonlinear conjugate gradient methods and Polyak’s heavy-ball approach, and is thus of high practical and theoretical interest in large-scale nonlinear optimization. We propose a general framework where the scalars of the linear combination defining the search direction are computed simultaneously by minimizing the approximate quadratic model in the 2 dimensional subspace. This strategy allows us to define a class of gradient methods with momentum enjoying global convergence guarantees and an optimal worst-case complexity bound in the nonconvex setting. Differently than all related works in the literature, the convergence conditions are stated in terms of the Hessian matrix of the bi-dimensional quadratic model. To the best of our knowledge, these results are novel to the literature. Moreover, extensive computational experiments show that the gradient methods with momentum here presented outperform classical conjugate gradient methods and are (at least) competitive with the state-of-art method for unconstrained optimization, i.e, L-BFGS method.
@@ 1 Introduction
To this aim, we draw inspiration from the idea presented in [11], where the search direction is computed by minimizing the approximate quadratic model in the 2 dimensional subspace spanned by the current gradient and the last search direction. According to this approach, the scalars $\alpha_{k}$ and $\beta_{k}$ are not prefixed, but rather they are simultaneously determined by a bidimensional search. We define a general framework of gradient methods with momentum based on a simple Armijo-type line search, and we prove global convergence results under the first-order smoothness assumptions only. We also derive specific algorithms with momentum from the general framework. Furthermore, we provide complexity results, proving for the proposed gradient method with momentum the worst-case complexity bound of ${\cal O}(\epsilon^{-2})$, which is optimal for first order algorithms in the nonconvex setting [18]. Up to our knowledge, the presented algorithm is the first framework of gradient methods with momentum having in the nonconvex case both theoretical convergence properties and a complexity bound.
Extensive computational experiments show that the gradient methods with momentum here presented outperform conjugate gradient methods and are (at least) competitive with the state-of-art method for unconstrained optimization, i.e, L-BFGS method.
@@ 2 The main idea
where $B_{k}$ is a suitable symmetric positive definite matrix. By properly choosing $B_{k}$, we retrieve standard methods; in particular:
@@
$f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (10)
@@
$f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}H_{11}&H_{12}\\ H_{12}&H_{22}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (11)
@@
From a theoretical point of view, the following issues must me considered:
@@
    we shall analyze under which conditions the obtained search direction $d_{k}$, coupled with line search techniques, allows to prove global convergence properties and complexity results for the iterative scheme (9).
@@
2.1 Related Works
We now briefly analyze some important works directly related to our approach.
Subspace minimization CG
Subspace minimization conjugate gradient (SMCG) methods, for which we refer the reader, for example, to the very recent paper [8], consider the two-dimensional subproblem (4)-(5) by assuming that the $n\times n$ matrix $B_{k}$ satisfies the secant equation
@@ 2.1 Related Works
The issue deserving attention becomes that of suitably managing the term $g_{k}^{T}B_{k}g_{k}=\rho_{k}$. Several strategies have been proposed in connection with this issue. Some global convergence results are stated, while complexity results are not known. However, as clearly written in [8], “A question is naturally to be asked: can one develop an efficient SMCG method without determining the parameter $\rho_{k}$?”. By our work, we give a positive answer to this question.
@@
The framework discussed in [19] considers (7)-(9) in a more general setting, i.e., by assuming that $P_{k}$ is a general $n\times m_{k}$ matrix, with $m_{k}\geq 1$, containing at least a search direction which satisfies an angle condition. The columns of $P_{k}$ are required to be linearly independent. Convergence results are stated by assuming that $\{B_{k}\}$ is a sequence of uniformly positive definite symmetric matrices bounded above. Then, gradient methods with momentum fits this framework provided that $g_{k}$ and $s_{k}$ are linearly independent. The convergence analysis relies on the properties of the $n\times n$ matrix $B_{k}$.
@@
Our main theoretical contribution, with respect to all the works on SMCG methods and to [19], concerns the definition of conditions on the $2\times 2$ matrix $H_{k}$ sufficient to ensure that the sequence of search directions $\{d_{k}\}$ is gradient-related and hence to state convergence and complexity results for the proposed gradient methods with momentum. As regards the contribution with respect to [21], we do not state assumptions on the Hessian matrix of the objective function: we only require that the gradient is Lipschitz continuous. Thus, we attain the same complexity result under much weaker assumptions. We remark that the theoretical focus on the low-dimensional matrix $H_{k}$ is fundamental to design convergent algorithms for large-scale optimization regardless of the $n\times n$ matrix $B_{k}$. This issue yields relevant implications, deeply analyzed and discussed later, from theoretical, algorithmic and computational point of views.
@@ 4 Properties of the search directions
where $\eta_{k}$ is the stepsize computed by an Armjio-type line search, is globally convergent to stationary points, Moreover, under this assumption the algorithm can be proved to have a worst case iteration and function evaluations complexity bound of ${\cal O}(\epsilon^{-2})$ to reach a solution $\bar{x}$ with $\|\nabla f(\bar{x})\|\leq\epsilon$. Note that this bound is actually tight for first-order methods under standard first-order smoothness assumptions.
@@ 4.1 Conditions on ${B_{k}}$
$g_{k}^{T}d_{k}\leq-\frac{\eta_{1}}{\eta_{2}^{2}}\|g_{k}\|^{2}$ (15)
$\frac{1}{\eta_{2}}\|g_{k}\|\leq\|d_{k}\|\leq\frac{2}{\eta_{1}}\|g_{k}\|.$ (16)
@@ 4.2 Conditions on $H_{k}$
By the next proposition, we finally state suitable conditions on the sequence of matrices $\{H_{k}\}$, that also take into account the sequences of vectors involved, i.e., $\{g_{k}\}$ and $\{s_{k}\}$. These conditions are sufficient to ensure that the obtained sequence of search directions $\{d_{k}\}$ is gradient-related.
@@ 5 Algorithmic model
In this section we propose an algorithmic framework for gradient methods with momentum which exploits the theoretical analysis carried out in the previous sections. The idea is to define a first-order algorithm with both strong theoretical guarantees in the nonconvex setting and the possibility of computationally exploiting eventual second-order information on the minimization problem. Before formally presenting the algorithm, we summarize the key steps:
@@
1 Input: $x_{0}\in\mathbb{R}^{n}$, $\gamma\in(0,1)$, $\delta\in(0,1)$, $c_{1}>0$, $c_{2}>0$.
2 Set $k\leftarrow 0$
3 while $\|\nabla f(x_{k})\|>0$ do
       /* Compute the search direction */
4       Set gr_dir_found $\leftarrow$ False
5       Define a $2\times 2$ symmetric matrix $H_{k}$
6       if problem (37) admits solution $[\alpha_{k}\;\;\beta_{k}]^{T}$ then
7             Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
8              if $g_{k}^{T}d_{k}\leq-c_{1}\|g_{k}\|^{2}$ and $\|d_{k}\|\leq c_{2}\|g_{k}\|$  then
9                    Set gr_dir_found $\leftarrow$ True
10
11      if gr_dir_found $=$ False then
12             Define a new $2\times 2$ symmetric matrix $H_{k}$ satisfying condition (28)
13              Compute $\alpha_{k}$ and $\beta_{k}$ by solving (37)
14             Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
15
      /* Perform Armijo line search along $d_{k}$/
16
17      Set $\eta\leftarrow 1$
18       while $f(x_{k}+\eta d_{k})>f(x_{k})+\gamma\eta d_{k}^{\top}\nabla f(x_{k})$ do
19              Set $\eta\leftarrow\delta\eta$
20
21      Set $\eta_{k}\leftarrow\eta$, $x_{k+1}\leftarrow x_{k}+\eta_{k}d_{k}$
22       Set $k\leftarrow k+1$
@@
From the steps the algorithm and Proposition 5 we have that $\{d_{k}\}$ is a sequence of gradient-related directions. Since the Armijo line search is employed within Algorithm 1, the results follow from [24, Proposition 1.2.1] and by [23]. ∎
We now focus on the issue of modifying a given $2\times 2$ matrix $H_{k}$ in order to satisfy condition (28) as required at step 1 of Algorithm 1. Suppose that a symmetric $H_{k}$ matrix has been defined at step 1, but either problem problem (37) does not admit solution or the test at step 1 is not satisfied, i.e., the value of gr_dir_found remains False. Then, step 1 must be performed and a new matrix $H_{k}$ must be constructed modifying as least as possible the given matrix defined at step 1.
Let us denote by $H_{k}^{0}$ the matrix defined at step 1 and by $H_{k}$ the new matrix defined at step 1. We can proceed as follows:
    Let $\hat{H}_{k}$ be obtained by a modified Cholesky factorization [24] applied to ${D_{k}^{-1}}H_{k}^{0}D_{k}^{-1}$, where
    $D_{k}=\begin{bmatrix}\|g_{k}\|&0\\ 0&\|s_{k}\|\end{bmatrix};$
    Set $H_{k}=D_{k}\hat{H}_{k}D_{k}$, i.e.,
@@
We conclude this section by stating a theoretical result showing that, under suitable assumptions on the objective function and a choice of the matrix $H_{k}$ actually related to the Hessian $\nabla^{2}f(x_{k})$, steps 1-1 are never executed for $k$ sufficiently large, that is, the test at step 1 is always satisfied.
@@
is the matrix defined at step 1, and let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Assume that $\{x_{k}\}$ converges to $x^{*}$, where $\nabla f(x^{*})=0$ and the Hessian matrix $\nabla^{2}f(x^{*})$ is positive definite. Furthermore, assume that the constants of the test at step 1 are chosen in a such a way that:
@@
where $\theta\in(0,1)$. Then, for $k$ sufficiently large the test at step 1 is satisfied.
@@ 6.1 Approximating Hessian-vector products by finite differences of gradients
In line with the strategy employed in [20, 21], we can draw inspiration by the Truncated Newton methods approach [25, 26], where the explicit management of the Hessian matrix $\nabla^{2}f(x_{k})$ is not required, but rather the Hessian-vector product $\nabla^{2}f(x_{k})d$ is directly handled, with $d\in R^{n}$.
@@
$H_{k}=\begin{bmatrix}g_{k}^{T}B_{k}g_{k}&-g_{k}^{T}B_{k}s_{k}\\ -g_{k}^{T}B_{k}s_{k}&s_{k}^{T}B_{k}s_{k}\end{bmatrix},$

### v3

@@ Abstract
In this work, we consider smooth unconstrained optimization problems and we deal with the class of gradient methods with momentum, i.e., descent algorithms where the search direction is defined as a linear combination of the current gradient and the preceding search direction. This family of algorithms includes nonlinear conjugate gradient methods and Polyak’s heavy-ball approach, and is thus of high practical and theoretical interest in large-scale nonlinear optimization. We propose a general framework where the scalars of the linear combination defining the search direction are computed simultaneously by minimizing the approximate quadratic model in the 2 dimensional subspace. This strategy allows us to define a class of gradient methods with momentum enjoying global convergence guarantees and an optimal worst-case complexity bound in the nonconvex setting. Differently than all related works in the literature, the convergence conditions are stated in terms of the Hessian matrix of the bi-dimensional quadratic model. To the best of our knowledge, these results are novel to the literature. Moreover, extensive computational experiments show that the gradient method with momentum here presented is a solid choice to tackle some classes of nonconvex unconstrained problems.
@@ 1 Introduction
To this aim, we draw inspiration from the idea presented in [11], where the search direction is computed by minimizing the approximate quadratic model in the 2 dimensional subspace spanned by the current gradient and the last search direction. According to this approach, the scalars $\alpha_{k}$ and $\beta_{k}$ are not prefixed, but rather they are simultaneously determined by a bidimensional search. We define a general framework of gradient methods with momentum based on a simple Armijo-type line search, and we prove global convergence results under the first-order smoothness assumptions only. We also derive specific algorithms with momentum from the general framework. Furthermore, we provide complexity results, proving for the proposed gradient method with momentum the worst-case complexity bound of ${\cal O}(\epsilon^{-2})$, which is optimal for first order algorithms in the nonconvex setting [18].
Up to our knowledge, the theoretical results presented here are novel in the literature. Known convergence and complexity properties for similar algorithmic frameworks are indeed based on assumptions that are stronger than ours or, as an alternative, can only be enforced in practice resorting to restarting techniques. Our analysis here does not make hypotheses concerning second order information for the objective nor linear independence between gradient and momentum directions; in addition, the conditions resulting from our analysis can be rigorously handled directly in the subspace, which is a crucial aspect from the computational perspective.
Extensive numerical experiments show that the gradient methods with momentum here presented outperform popular implementations of the conjugate gradient method and are (at least) competitive with a widespread limited-memory Quasi-Newton method like L-BFGS algorithm [19]. Moreover, while the conjugate gradient based procedure implemented in cg_descent software [20, 21] proves to still remain superior, the proposed method seems comparable, with some upsides, in the particular, relevant class of multi-layer perceptron training problems [22, 23].
@@ 2 The main idea
where $B_{k}$ is a suitable symmetric positive definite matrix. By properly choosing $B_{k}$, we retrieve standard methods, in particular:
@@
$\psi_{k}(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (10)
@@
$\psi_{k}(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}H_{11}&H_{12}\\ H_{12}&H_{22}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (11)
@@
From a theoretical point of view, the following issues must be considered:
@@
    we shall analyze under which conditions the obtained search direction $d_{k}$, coupled with line search techniques, allow to prove global convergence properties and complexity results for the iterative scheme (9).
@@
2.1 Related works and open problems
Before turning to the theoretical analysis, we shall briefly discuss some important works directly related to the approach addressed in this paper.
Subspace minimization conjugate gradient (SMCG) methods
The class of SMCG type algorithms, for which we refer the reader, for example, to the very recent paper [8], consider the two-dimensional subproblem (4)-(5) by assuming that the $n\times n$ matrix $B_{k}$ satisfies the secant equation
@@ 2.1 Related works and open problems
The issue deserving attention becomes that of suitably managing the term $g_{k}^{T}B_{k}g_{k}=\rho_{k}$. Several strategies have been proposed in connection with this issue. Some global convergence results are stated, while complexity results are not known. However, as clearly written in [8], “A question is naturally to be asked: can one develop an efficient SMCG method without determining the parameter $\rho_{k}$?”. In other words, is it possible to define a convergent SMCG method without the need of handling the full matrix $B_{k}$? By our work, we give a positive answer to this question.
@@
The framework discussed in [24] considers (7)-(9) in a more general setting, i.e., by assuming that $P_{k}$ is a general $n\times m_{k}$ matrix, with $m_{k}\geq 1$, containing at least a search direction which satisfies an angle condition. The columns of $P_{k}$ are required to be linearly independent. Convergence results are stated by assuming that $\{B_{k}\}$ is a sequence of uniformly positive definite symmetric matrices bounded above. However, while the convergence analysis relies on the properties of the $n\times n$ matrix $B_{k}$, it is crucial, for computational performance, to only deal with the small matrix $H_{k}$.
@@
Our main theoretical contribution with respect to all the works on SMCG methods and to [24], thus concerns the definition of conditions on the $2\times 2$ matrix $H_{k}$ sufficient to ensure that the sequence of search directions $\{d_{k}\}$ is gradient-related and hence to state convergence and complexity results for the proposed gradient methods with momentum. Note that working with only the $2\times 2$ matrix, regardless of the $n\times n$ matrix $B_{k}$, is fundamental to make these algorithmic frameworks computationally viable at large scale; our result on this issue thus enlightens how it can be done in a theoretically sound manner.
As regards the contribution with respect to [26], we do not state assumptions on the Hessian matrix of the objective function: we only require that the gradient is Lipschitz continuous. Thus, we attain the same complexity result under much weaker assumptions.
@@ 4 Properties of the search directions
where $\eta_{k}$ is the stepsize computed by an Armjio-type line search, is globally convergent to stationary points. Moreover, under this assumption the algorithm can be proved to have a worst case iteration and function evaluations complexity bound of ${\cal O}(\epsilon^{-2})$ to reach a solution $\bar{x}$ with $\|\nabla f(\bar{x})\|\leq\epsilon$. Note that this bound is actually tight for first-order methods under standard first-order smoothness assumptions.
@@ 4.1 Conditions on ${B_{k}}$
$g_{k}^{T}d_{k}\leq-\frac{\eta_{1}}{\eta_{2}^{2}}\|g_{k}\|^{2}$
$\frac{1}{\eta_{2}}\|g_{k}\|\leq\|d_{k}\|\leq\frac{2}{\eta_{1}}\|g_{k}\|.$
@@ 4.2 Conditions on $H_{k}$
By the next proposition, we finally state suitable conditions on the sequence of matrices $\{H_{k}\}$, that also take into account the sequences of vectors involved, i.e., $\{g_{k}\}$ and $\{s_{k}\}$. These conditions represent the main theoretical contribution of the paper, as they are sufficient to ensure that the obtained sequence of search directions $\{d_{k}\}$ is gradient-related.
@@
We will show later in this work how the condition can practically be enforced in computational scenarios.
@@ 5 Algorithmic model
In this section we present an algorithmic framework for gradient methods with momentum which exploits the theoretical analysis carried out in the previous sections. The idea is to define a first-order algorithm with both strong theoretical guarantees in the nonconvex setting and the possibility of computationally exploiting eventual second-order information on the minimization problem. Before formally presenting the algorithm, we summarize the key steps:
@@
1:Input: $x_{0}\in\mathbb{R}^{n}$, $\gamma\in(0,1)$, $\delta\in(0,1)$, $c_{1}>0$, $c_{2}>0$.
2:Set $k\leftarrow 0$
3:while $\|\nabla f(x_{k})\|>0$ do
4:     /$\ast$/ ∗ Compute the search direction $\ast$/∗ /
5:     Set gr_dir_found $\leftarrow$ False
6:     Define a $2\times 2$ symmetric matrix $H_{k}$
7:     if problem (35) admits solution $[\alpha_{k}\;\;\beta_{k}]^{T}$ then
8:         Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
9:         if $g_{k}^{T}d_{k}\leq-c_{1}\|g_{k}\|^{2}$ and $\|d_{k}\|\leq c_{2}\|g_{k}\|$  then
10:              Set gr_dir_found $\leftarrow$ True
11:     if gr_dir_found $=$ False then
12:         Define a new $2\times 2$ symmetric matrix $H_{k}$ satisfying condition (26)
13:         Compute $\alpha_{k}$ and $\beta_{k}$ by solving (35)
14:         Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
15:     /$\ast$/ ∗ Perform Armijo line search along $d_{k}$ $\ast$/∗ /
16:     Set $\eta\leftarrow 1$
17:     while $f(x_{k}+\eta d_{k})>f(x_{k})+\gamma\eta d_{k}^{\top}\nabla f(x_{k})$ do
18:         Set $\eta\leftarrow\delta\eta$
19:     Set $\eta_{k}\leftarrow\eta$, $x_{k+1}\leftarrow x_{k}+\eta_{k}d_{k}$
20:     Set $k\leftarrow k+1$
@@
From the steps the algorithm and Proposition 5 we have that $\{d_{k}\}$ is a sequence of gradient-related directions. Since the Armijo line search is employed within Algorithm 1, the results follow from [29, Proposition 1.2.1] and by [28]. ∎
We now focus on the issue of modifying a given $2\times 2$ matrix $H_{k}$ in order to satisfy condition (26) as required at step 12 of Algorithm 1. Suppose that a symmetric $H_{k}$ matrix has been defined at step 6, but either problem (35) does not admit solution or the test at step 9 is not satisfied, i.e., the value of gr_dir_found remains False. Then, step 12 must be performed and a new matrix $H_{k}$ must be constructed modifying as least as possible the given matrix defined at step 6.
Let us denote by $H_{k}^{0}$ the matrix defined at step 6 and by $H_{k}$ the new matrix defined at step 12. We can proceed as follows:
    A matrix $\hat{H}_{k}$ can be obtained by a modified Cholesky factorization [29] applied to $H_{k}^{0}$
    Letting
    $D_{k}=\begin{bmatrix}\|g_{k}\|&0\\ 0&\|s_{k}\|\end{bmatrix}$
    we can set $H_{k}=D_{k}\hat{H}_{k}D_{k}$, i.e.,
@@
We conclude this section by stating a theoretical result showing that, under suitable assumptions on the objective function and a choice of the matrix $H_{k}$ actually related to the Hessian $\nabla^{2}f(x_{k})$, steps 12-14 are never executed for $k$ sufficiently large, that is, the test at step 9 is always satisfied.
@@
is the matrix defined at step 6, and let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Assume that $\{x_{k}\}$ converges to $x^{*}$, where $\nabla f(x^{*})=0$ and the Hessian matrix $\nabla^{2}f(x^{*})$ is positive definite. Furthermore, assume that the constants of the test at step 9 are chosen in a such a way that:
@@
where $\theta\in(0,1)$. Then, for $k$ sufficiently large the test at step 9 is satisfied.
@@ 6.1 Approximating Hessian-vector products by finite differences of gradients
In line with the strategy employed in [25, 26], we can draw inspiration by the strategy employed in Truncated Newton methods [30, 31], where the explicit management of the Hessian matrix $\nabla^{2}f(x_{k})$ is not required, but rather the Hessian-vector product $\nabla^{2}f(x_{k})d$ is directly handled, with $d\in R^{n}$.
@@
$H_{k}=\begin{bmatrix}g_{k}^{T}B_{k}g_{k}&-g_{k}^{T}B_{k}s_{k}\\ -g_{k}^{T}B_{k}s_{k}&s_{k}^{T}B_{k}s_{k}\end{bmatrix}.$